OpenStudy (anonymous):

A circle has a perimeter of 35 feet. What is its diameter?

OpenStudy (nurali):

C = pi * d where pi = 3.14 35 = 3.14*d d=35/3.14 d = 11.1 feet

OpenStudy (goformit100):



OpenStudy (cggurumanjunath):